# ARISE – Adaptive Refinement and Iterative Scenario Engineering

### v1

@@ Abstract
The effectiveness of collision-free trajectory planners depends on the quality and diversity of training data, especially for rare scenarios. A widely used approach to improve dataset diversity involves generating realistic synthetic traffic scenarios. However, producing such scenarios remains difficult due to the precision required when scripting them manually or generating them in a single pass. Natural language offers a flexible way to describe scenarios, but existing text-to-simulation pipelines often rely on static snippet retrieval, limited grammar, single-pass decoding, or lack robust executability checks. Moreover, they depend heavily on constrained LLM prompting with minimal post-processing. To address these limitations, we introduce ARISE — Adaptive Refinement and Iterative Scenario Engineering, a multi-stage tool that converts natural language prompts into executable Scenic scripts through iterative LLM-guided refinement. After each generation, ARISE tests script executability in simulation software, feeding structured diagnostics back to the LLM until both syntactic and functional requirements are met. This process significantly reduces the need for manual intervention. Through extensive evaluation, ARISE outperforms the baseline in generating semantically accurate and executable traffic scenarios with greater reliability and robustness.
@@ I Introduction
Autonomous driving (AD) systems require thorough testing across a wide spectrum of traffic scenarios, including out-of-distribution scenarios, which are either rarely seen in the real-world, or impractical to capture, due to the underlying costs, and danger to be staged in physical experiments. Thus, simulation-based traffic scenario generation has emerged as a practical and cost-efficient alternative to real-world testing. The comprehensive and cost-efficient, safety-critical scenarios can be potentially created on demand using the domain-specific tools like Scenic [16] and OpenSCENARIO [1], and thereby expose autonomous vehicle (AV) models to unusual and hazardous situations and conditions.
Large Language Models (LLMs) have recently gained researchers’ attention for transforming natural language descriptions into executable simulated traffic scenarios. Using the rich knowledge encapsulated in them, LLMs allow for the direct translation of the plain text describing a scenario into the simulation specific code. ChatScene [38] introduced an LLM-based tool that generates driving scenarios in Scenic format, executable within the CARLA simulator [^2] https://carla.org [10], by integrating a structured knowledge base with LLM-generated responses to construct runnable scripts. Numerous tools (described in more detail in the Related Work section) leverage LLMs to transform free-form text descriptions (e.g. traffic accident reports, short prompts) into the Scenic scripts. These works demonstrate the advantages of using LLMs to automate traffic scenario generation.
@@
Although there is a clear trend toward future scenario generation pipelines that rely on human-language-based precision, the challenge of reliably translating natural language into executable simulation scripts remains unresolved. Compilation, simulation and runtime errors might render a resulting script unusable. Existing works tend to not test the executability of resulting scripts, requiring manual corrections in case of errors. This undermines the intended automation of such pipelines and highlights the need for a more robust and adaptive solution capable of producing reliable and refined results.
To tackle these challenges, we present ARISE: Adaptive Refinement and Iterative Scenario Engineering, an advanced tool for simulated traffic scenario generation from natural language descriptions (cf. Figure 1 for example). ARISE directly extends the open-source ChatScene codebase by integrating an important Test-and-Repair Loop (TRL) into the generation pipeline. Similar to the ChatScene, ARISE initially combines the knowledge base, prompting and LLM to draft a candidate Scenic script, which is then automatically compiles and tests in the CARLA simulator. When errors are detected, ARISE automatically initiates an iterative repair process, significantly minimizing the need for user intervention. Feeding the diagnostic feedback (e.g. error stack traces, debug output) from the previous attempts back to the LLM with a structured prompt, the TRL seeks to produce revised and valid output. Thus, ARISE transforms the single-pass generation into a dynamic dialogue between the generator and the LLM. Leveraging the continuous advancement mechanism, ARISE offers improved success rates and practical robustness when compared to earlier works. In our experiments, ARISE achieves higher Execution Success Rate (ESR) (1) - the ratio of the resulting traffic scenario scripts that execute without error without user intervention, to the total amount of generated scripts. Additionally, we use Repair Convergence Rate (RCR) (2) that indicates how quickly the repair loop converges to a valid scenario, and Semantical Conformity Score (SCS) (3), which provides insight in the semantical correspondence between the initial scenario descriptions and resulting Scenic scripts. The experiments demonstrate very good results, proving the efficiency of the proposed advancements.
@@
  • Improved performance, where ARISE outperforms prior approach in terms of execution success and repair convergence rates (ESR and RCR respectively).
  • Robustness and contextual consistency, where we demonstrate that ARISE delivers results that are not only syntactically correct and executable, but also are semantically conform with initial descriptions.
@@ II-B LLM-empowered Frameworks
Large Language Models (LLMs) have recently shown strong potential in traffic scenario generation, with several systems exploring different strategies to combine language understanding with scene synthesis. Some approaches guide generative models to create detailed traffic scenes or visual assets. For instance, CTG++ [41] introduces a scene-level LLM-guided conditional diffusion model, blending the strengths of diffusion processes with textual conditioning to generate diverse traffic scenes. Similarly, ChatSim [34] employs LLMs in combination with image and video generators to create photo-realistic assets and scene videos, aiming at high-fidelity visual outputs.
@@
To enhance controllability and scene realism, retrieval-augmented and iterative refinement approaches have also emerged. OmniTester [22] combines an LLM with the SUMO road network generator and Retrieval-Augmented Generation (RAG) [5], enabling generation of both dynamic elements and static infrastructure from user prompts. Additionally, an LLM-based evaluator is used to iteratively refine outputs, focusing on background vehicles behavior.
@@ II-C LLM-based Scenario Generation Using Scenic
Recent works have focused on leveraging large language models for automatic generation of synthetic traffic scenarios from textual descriptions. ChatScene [38] leverages a combination of an LLM with a knowledge base containing domain examples to produce scripts encoding traffic scenarios in single pass. Given a user prompt input, ChatScene decomposes the scenario into a set of key components: relative spawn positions, adversarial behavior and road geometry. It then retrieves the matching Scenic code snippets from the database, and assembles them into a full script that can be potentially executed in the CARLA simulation software.
@@ III-A Scenario Prompts
In line with the baseline, we use 40 prompts describing diverse safety-critical traffic scenarios (each describing interaction between ego vehicle and adversarial agent, with various levels of other details), grouped into eight categories, depicted in the header of Table I. Each scenario is defined through a natural language description. A sample of a scenario prompt (starting point) is depicted on Figure 2 (module A, upper left).
@@ III-B Module A: Extraction
The extraction stage follows the approach of ChatScene, with the primary distinction being the extension of semantic sub-categories within scenario descriptions. First, ARISE processes natural language descriptions input of traffic scenarios. Second, an LLM is prompted using a few-shot template that demonstrates how to identify various semantic components. Third, the LLM model then extracts key structural elements.
@@ III-E Module D: Test-and-Repair Loop (TRL)
If compilation and execution tests fail, then the script is passed on to the Test-and-Repair Loop (TRL) (cf. Figure 2, module D) for further postprocessing. ARISE differs significantly from the baseline by introducing the automated, iterative TR. This module constitutes the core novelty of our approach. While the baseline follows a single-pass pipeline, producing Scenic scripts without validation and requiring manual debugging in case of errors, ARISE aims to enhance automation by iteratively repairing scripts to produce fully functional outputs with no user involvement.
In each iteration, structured diagnostic feedback, including error messages, stack traces, and debug information, is passed to an LLM via a carefully crafted automated few-shot prompt. This prompt includes the original scenario description, the most recent script version, and detailed instructions along with selected examples and code snippets specific to the TRL stage. The LLM is tasked with identifying and correcting the errors of the script based on this feedback. The revised script is then recompiled and re-tested in simulation. This process repeats until either the script passes both stages successfully or a predefined maximum number of iterations is reached.
If the script passes both compilation and simulation steps, it is saved, and relevant statistics are recorded. If all repair attempts fail, the final version of the script is still saved, allowing for manual correction. Furthermore, ARISE then proceeds to the next execution run or next scenario.
Importantly, ARISE transforms the generation pipeline into an autonomous and adaptive debugging agent, acting as a mediator between our tool, the simulator and the LLM. This results in significantly improved robustness and execution success rates compared to baseline. For demonstration purposes, in Figure 3 we present a sequence of frames illustrating key events from a safety-critical scenario rendered in the CARLA.
@@ IV Evaluations
We conducted experiments using 40 natural language scenario descriptions, consistent with the baseline setup. These descriptions are grouped into 8 categories (or scenario types), with 5 distinct prompts per category. Given the inherent non-determinism of LLMs, and to improve the robustness of the evaluation, we executed each method multiple times per scenario (typically 50, and in some cases, 100 runs - see Table I). All experiments were carried out using OpenAI’s GPT-4o model. The maximum number of iterations for TRL was set to 10, and the Scenic simulation environment was allowed up to 15 spawn attempts per script. Since baseline bypasses LLM-based refinement when the top-k retrieval parameter is set to 1 (default), we set top-k = 2 for all evaluated methods to ensure consistency and comparability.
@@ IV-A Metrics
Since ARISE aims to reduce the need for manual intervention in producing executable traffic scenario scripts, we introduce two evaluation metrics: Execution Success Rate (ESR) and Repair Convergence Rate (RCR). ESR measures the proportion of Scenic scripts that are not only syntactically valid and compilable, but also are executable in CARLA. The ESR is computed as:
@@ IV-B Quantitative Evaluations
ARISE significantly outperforms (average ESR of 78%) both the baseline (cf. Table I), which often failed to produce valid simulations in several scenario categories, resulting in ESR values close to zero (average of 1.4%), and the baseline that we extended with TRL (modified, average of 20.5%). Note, RCR metric is not applicable to the baseline since it apperates in a single pass script generation mode. We have implemented the latter as an ablation to highlight the isolated impact of the individual improvements that we have introduced in ARISE. These results highlight the effectiveness of combining TRL with an expanded knowledge base and fine-grained semantic prompting, significantly reducing the need for manual involvement to produce syntactically correct and executable scripts across all evaluated traffic categories. Further improvements are observed when lowering the LLM temperature to 0.3, which yields better ESR and RCR scores compared to both the baseline and ARISE with the default temperature of 1.0. Lowering the temperature increases determinism, while still allowing for limited output diversity. A more comprehensive exploration of temperature and other fine-tuning parameters is left for future work. For completeness, we additionally report ChatScene’s one-shot (pre-TRL) results to benchmark the pure baseline. While the overall ESR values are nearly identical (0.014 for ChatScene and 0.012 for ChatScene + TRL), performance varied across scenario categories. These differences can be attributed to the inherent non-determinism of LLM outputs. RCR results indicate steady improvement across the methods. Notably, ARISE’s overall success rate was over an order of magnitude higher than the baseline, and significantly higher than the baseline with TRL. This improvement supports our claim of the effectiveness of the proposed combined approach with TRL, more fine-granular extraction prompts, and an extended (and revised) knowledge base. Moreover, RCR numbers show that ARISE, on average, requires a relatively small number of repair cycles per scenario. Most baseline scenarios were automatically converted into functional Scenic scripts, advancing our goal of a fully automated traffic scenario generation pipeline and its importance for community.
@@ IV-C Qualitative Evaluations
Inspired by OmniTester’s LLM-based evaluator, we introduce our own module to assess the semantic correspondence between scenario descriptions and generated Scenic scripts using Equation 3. The evaluator, implemented with Google Gemini 2.5 Flash, operates in two stages: it is first primed with a condensed Scenic reference and curated examples, then evaluates description–script pairs using a structured few-shot prompt with seven scoring criteria (adversarial type, behavior, geometry, weather, elements, spawn, and requirements). We conduct two evaluations—inter-scenario/inter-LLM SCS comparison and scoring consistency analysis (cf. Table II).
Semantical Evaluation 1. In the first experiment, we selected first 10 generated scripts per scenario type for each LLM, using the same criteria as in the inter-LLM ESR and RCR comparison (i.e., the first scenario from each category) and evaluated them using our LLM-based evaluator. The key results are reported in Table II. The results demonstrate inter-LLM inter-scenario type average SCS varying between 84.9% (Gemini 2.0 Flash) and 88.2% (GPT-4o, temperature=0.3), while the lowest inter-scenario scores lies at 50% (GPT-4o, temperature 1.0, Lane Changing; Gemini 2.0 Flash, temperature=0.3, Red-Light Running). Showing such a high average SCS further improves our confidence in ARISE’s ability to preserve the semantical conformity between the initial scenario description and the resulting Scenic scripts.
Semantical Evaluation 2. To assess the soundness of our approach, we analyze the standard deviation of SCS in two settings (cf. Table II): repeated scoring of the same script (10 times each across 8 scenario types), and scoring across 8 groups of 10 different scripts (one group per scenario type). This allows to evaluate the consistency and reliability of the LLM-based evaluator, expecting similar scores for the same and semantically similar scripts. We exclude other LLMs here, as the goal is to evaluate the stability of the evaluator itself, and including additional models would provide limited added value. Table III shows a low average standard deviation of 3.75 percentage points when the same script is evaluated repeatedly across scenario categories. In contrast, Table II presents higher standard deviations in the inter-scenario, inter-LLM comparison involving different scripts (10 per scenario), which is expected due to script variability. These results indicate that the LLM-based semantic evaluator produces consistent scores, supporting the reliability of the scoring approach.

### v2

@@ Abstract
The effectiveness of collision-free trajectory planners depends on the quality and diversity of training data, especially for rare scenarios. A widely used approach to improve dataset diversity involves generating realistic synthetic traffic scenarios. However, producing such scenarios remains difficult due to the precision required when scripting them manually or generating them in a single pass. Natural language offers a flexible way to describe scenarios, but existing text-to-simulation pipelines often rely on static snippet retrieval, limited grammar, single-pass decoding, or lack robust executability checks. Moreover, they depend heavily on constrained LLM prompting with minimal post-processing. To address these limitations, we introduce ARISE 2Adaptive Refinement and Iterative Scenario Engineering, a multi-stage tool that converts natural language prompts into executable Scenic scripts through iterative LLM-guided refinement. After each generation, ARISE tests script executability in simulation software, feeding structured diagnostics back to the LLM until both syntactic and functional requirements are met. This process significantly reduces the need for manual intervention. Through extensive evaluation, ARISE outperforms the baseline in generating semantically accurate and executable traffic scenarios with greater reliability and robustness.
@@ I Introduction
Autonomous driving (AD) systems require thorough testing across a wide spectrum of traffic scenarios, including out-of-distribution scenarios, which are either rarely seen in the real-world, or impractical to capture, due to the underlying costs, and danger to be staged in physical experiments. Thus, simulation-based traffic scenario generation has emerged as a practical and cost-efficient alternative to real-world testing. The comprehensive and cost-efficient safety-critical scenarios can potentially be created on demand using the domain-specific tools like Scenic [16] and OpenSCENARIO [1], and thereby expose autonomous vehicle (AV) models to unusual and hazardous situations and conditions.
Large Language Models (LLMs) have recently gained researchers’ attention for transforming natural language descriptions into executable simulated traffic scenarios. Using the rich knowledge encapsulated in them, LLMs allow for the direct translation of the plain text describing a scenario into the simulation specific code. ChatScene [38] introduced an LLM-based tool that generates driving scenarios in Scenic format, executable within the CARLA simulator (https://carla.org[10], by integrating a structured knowledge base with LLM-generated responses to construct runnable scripts. Numerous tools (described in more detail in the Related Work section) leverage LLMs to transform free-form text descriptions (e.g. traffic accident reports, short prompts) into the Scenic scripts. These works demonstrate the advantages of using LLMs to automate traffic scenario generation.
@@
Although there is a clear trend toward future scenario generation pipelines that rely on human-language-based precision, the challenge of reliably translating natural language into executable simulation scripts remains unresolved. Compilation, simulation and runtime errors might render a resulting script unusable. Existing works tend not to test the executability of resulting scripts, requiring manual corrections in case of errors. This undermines the intended automation of such pipelines and highlights the need for a more robust and adaptive solution capable of producing reliable and refined results.
To tackle these challenges, we present ARISE: Adaptive Refinement and Iterative Scenario Engineering, an advanced tool for simulated traffic scenario generation from natural language descriptions (cf. Figure 1). ARISE directly extends the open-source ChatScene codebase by integrating an important Test-and-Repair Loop (TRL) into the generation pipeline. Similar to the ChatScene, ARISE initially combines the knowledge base, prompting and LLM to draft a candidate Scenic script, which is then automatically compiled and tested in the CARLA simulator. When errors are detected, ARISE automatically initiates an iterative repair process, significantly minimizing the need for user intervention. By feeding the diagnostic feedback (e.g. error stack traces, debug output) from the previous attempts back to the LLM with a structured prompt, the TRL seeks to produce revised and valid output. Thus, ARISE transforms the single-pass generation into a dynamic dialogue between the generator and the LLM. Leveraging the continuous advancement mechanism, ARISE offers improved success rates and practical robustness when compared to earlier works. In our experiments, ARISE achieves higher Execution Success Rate (ESR) (1) - the ratio of the resulting traffic scenario scripts that execute without error and without user intervention, to the total number of generated scripts. Additionally, we use the Repair Convergence Rate (RCR) (2), which indicates how quickly the repair loop converges to a valid scenario, and the Semantical Conformity Score (SCS) (3), which provides insight into the semantical correspondence between the initial scenario descriptions and resulting Scenic scripts. The experiments demonstrate very good results, proving the efficiency of the proposed advancements.
@@
  • Improved performance, where ARISE outperforms the prior approach in terms of execution success and repair convergence rates (ESR and RCR respectively).
  • Robustness and contextual consistency, where we demonstrate that ARISE delivers results that are not only syntactically correct and executable, but are also semantically conform to the initial descriptions.
@@ II-B LLM-empowered Frameworks
LLMs have recently shown strong potential in traffic scenario generation, with several systems exploring different strategies to combine language understanding with scene synthesis. Some approaches guide generative models to create detailed traffic scenes or visual assets. For instance, CTG++ [41] introduces a scene-level LLM-guided conditional diffusion model, blending the strengths of diffusion processes with textual conditioning to generate diverse traffic scenes. Similarly, ChatSim [34] employs LLMs in combination with image and video generators to create photo-realistic assets and scene videos, aiming at high-fidelity visual outputs.
@@
To enhance controllability and scene realism, retrieval-augmented and iterative refinement approaches have also emerged. OmniTester [22] combines an LLM with the SUMO road network generator and Retrieval-Augmented Generation (RAG) [5], enabling generation of both dynamic elements and static infrastructure from user prompts. Additionally, an LLM-based evaluator is used to iteratively refine outputs, focusing on background vehicles’ behavior.
@@ II-C LLM-based Scenario Generation Using Scenic
Recent works have focused on leveraging LLMs for automatic generation of synthetic traffic scenarios from textual descriptions. ChatScene [38] leverages a combination of an LLM with a knowledge base containing domain examples to produce scripts encoding traffic scenarios in single pass. Given a user prompt input, ChatScene decomposes the scenario into a set of key components: relative spawn positions, adversarial behavior and road geometry. It then retrieves the matching Scenic code snippets from the database, and assembles them into a full script that can be potentially executed in the CARLA simulation software.
@@ III-A Scenario Prompts
In line with the baseline, we use 40 prompts describing diverse safety-critical traffic scenarios (each describing interaction between ego vehicle and adversarial agent, with various levels of other details), grouped into eight categories, depicted in the header of Table I. Each scenario is defined through a natural language description. A sample of a scenario prompt (starting point) is depicted in Figure 2 (module A, upper left).
@@ III-B Module A: Extraction
The extraction stage follows the approach of ChatScene, with the primary distinction being the extension of semantic sub-categories within scenario descriptions. First, ARISE processes natural language descriptions input of traffic scenarios. Second, an LLM is prompted using a few-shot template that demonstrates how to identify various semantic components. Third, the LLM then extracts key structural elements.
@@ III-E Module D: Test-and-Repair Loop (TRL)
If compilation and execution tests fail, then the script is passed on to the Test-and-Repair Loop (TRL) (cf. Figure 2, module D) for further post-processing. ARISE differs significantly from the baseline by introducing the automated, iterative TRL. This module constitutes the core novelty of our approach. While the baseline follows a single-pass pipeline, producing Scenic scripts without validation and requiring manual debugging in case of errors, ARISE aims to enhance automation by iteratively repairing scripts to produce fully functional outputs with no user involvement.
In each iteration, structured diagnostic feedback, including error messages, stack traces, and debug information, is passed to an LLM via a carefully crafted automated few-shot prompt. This prompt includes the original scenario description, the most recent script version, and detailed instructions along with selected examples and code snippets specific to the TRL stage. The LLM is tasked with identifying and correcting the errors of the script based on this feedback. The revised script is then recompiled and re-tested in simulation. This process repeats until either the script passes both stages successfully or a predefined maximum number of iterations (set by a user) is reached. If the script passes both compilation and simulation steps, it is saved, and relevant statistics are recorded. If all repair attempts fail, the final version of the script is still saved, allowing for manual correction. Furthermore, ARISE then proceeds to the next execution run or next scenario.
Importantly, ARISE transforms the generation pipeline into an autonomous and adaptive debugging agent, acting as a mediator between our tool, the simulator and the LLM. This results in significantly improved robustness and execution success rates compared to baseline. For demonstration purposes, in Figure 3, we present a sequence of frames illustrating key events from a safety-critical scenario rendered in the CARLA.
@@ IV Evaluations
We conducted experiments using 40 natural language scenario descriptions, consistent with the baseline setup. These descriptions are grouped into 8 categories (scenario types), with 5 distinct prompts per category. Given the inherent non-determinism of LLMs, and to improve the robustness of the evaluation, we executed each method multiple times per scenario (typically 50, and in some cases, 100 runs - see Table I). All experiments were carried out using OpenAI’s GPT-4o model. The maximum number of iterations for TRL was set to 10, and the Scenic simulation environment was allowed up to 15 spawn attempts per script. Since the baseline bypasses LLM-based refinement when the top-k retrieval parameter is set to 1 (default), we set top-k = 2 for all evaluated methods to ensure consistency and comparability.
@@ IV-A Metrics
Since ARISE aims to reduce the need for manual intervention in producing executable traffic scenario scripts, we introduce two evaluation metrics: Execution Success Rate (ESR) and Repair Convergence Rate (RCR). ESR measures the proportion of Scenic scripts that are not only syntactically valid and compilable, but also executable in CARLA. The ESR is computed as:
@@ IV-B Quantitative Evaluations
ARISE significantly outperforms (average ESR of 78%) both the baseline (cf. Table I), which often failed to produce valid simulations in several scenario categories, resulting in ESR values close to zero (average of 1.4%), and the baseline that we extended with TRL (modified, average of 20.5%). Note: the RCR metric is not applicable to the baseline since it operates in a single-pass script generation mode. We have implemented the latter as an ablation to highlight the isolated impact of the individual improvements that we have introduced in ARISE. These results highlight the effectiveness of combining TRL with an expanded knowledge base and fine-grained semantic prompting, significantly reducing the need for manual involvement to produce syntactically correct and executable scripts across all evaluated traffic categories. Further improvements are observed when lowering the LLM temperature to 0.3, which yields better ESR and RCR scores compared to both the baseline and ARISE with the default temperature of 1.0. Lowering the temperature increases determinism, while still allowing for limited output diversity. A more comprehensive exploration of temperature and other fine-tuning parameters is left for future work. For completeness, we additionally report ChatScene’s one-shot (pre-TRL) results to benchmark the pure baseline. While the overall ESR values are nearly identical (0.014 for ChatScene and 0.012 for ChatScene + TRL), performance varied across scenario categories. These differences can be attributed to the inherent non-determinism of LLM outputs. RCR results indicate steady improvement across the methods. Notably, ARISE’s overall success rate was over an order of magnitude higher than the baseline, and significantly higher than the baseline with TRL. This improvement supports our claim of the effectiveness of the proposed combined approach with TRL, more fine-granular extraction prompts, and an extended (and revised) knowledge base. Moreover, RCR numbers show that ARISE, on average, requires a relatively small number of repair cycles per scenario. Most baseline scenarios were automatically converted into functional Scenic scripts, advancing our goal of a fully automated traffic scenario generation pipeline and its importance for the community.
@@ IV-C Qualitative Evaluations
Inspired by OmniTester’s LLM-based evaluator, we introduce our own module to assess the semantic correspondence between scenario descriptions and generated Scenic scripts using Equation 3. The evaluator, implemented with Google Gemini 2.5 Flash, operates in two stages: it is first primed with a condensed Scenic reference and curated examples, then evaluates description–script pairs using a structured few-shot prompt with seven scoring criteria (adversarial type, behavior, geometry, weather, elements, spawn, and requirements). We conduct two evaluations: inter-scenario/inter-LLM SCS comparison and scoring consistency analysis (cf. Table II).
Semantical Evaluation 1. In the first experiment, we selected the first 10 generated scripts per scenario type for each LLM, using the same criteria as in the inter-LLM ESR and RCR comparison (i.e., the first scenario from each category) and evaluated them using our LLM-based evaluator. The key results are reported in Table II. The results demonstrate that the inter-LLM, inter-scenario-type average SCS varying between 84.9% (Gemini 2.0 Flash) and 88.2% (GPT-4o, temperature=0.3), while the lowest inter-scenario score lies at 50% (GPT-4o, temperature 1.0, Lane Changing; Gemini 2.0 Flash, temperature=0.3, Red-Light Running). Showing such a high average SCS further improves our confidence in ARISE’s ability to preserve the semantical conformity between the initial scenario description and the resulting Scenic scripts.
Semantical Evaluation 2. To assess the soundness of our approach, we analyze the standard deviation of SCS in two settings (cf. Table II): repeated scoring of the same script (10 times each across 8 scenario types), and scoring across 8 groups of 10 different scripts (one group per scenario type). This allows us to evaluate the consistency and reliability of the LLM-based evaluator, expecting similar scores for the same and semantically similar scripts. We exclude other LLMs here, as the goal is to evaluate the stability of the evaluator itself, and including additional models would provide limited added value. Table III shows a low average standard deviation of 3.75 percentage points when the same script is evaluated repeatedly across scenario categories. In contrast, Table II presents higher standard deviations in the inter-scenario, inter-LLM comparison involving different scripts (10 per scenario), which is expected due to script variability. These results indicate that the LLM-based semantic evaluator produces consistent scores, supporting the reliability of the scoring approach.